Together, a necklace and a braclet cost $192. Find the price of each if the necklace cost 3 times as much as the braclet.
Answer by shweta(56)

Let the cost of necklace be $N and bracelet be $B
Together they cost $192, so N+B= 192 ...(1)
given N = 3*B ...(2)
Substitute the value of 2 in 1
N+B=192
3B+B= 192
4B= 192
B= 192/4
B= 48
Cost of bracelet is $48
Put in the value of B in 1 to get the value of N
N+B= 192
N+48=192
N= 192-48
N= 144
Cost of necklace is $144